Session: Nov 2002 (session 3)
Wavelength: 21 cm
To give feedback on this experiment, click
Medicina (Franco Mantovani at Wed Nov 13 02:08:26 2002) Minor failures : RFI
Strong interferences at the observing frequency.
Onsala (Michael Lindqvist at Wed Nov 13 07:22:00 2002) Success :
No known problems.
Torun (K.Borkowski at Wed Nov 13 08:36:02 2002) Success : HA-E
Sources below telescope horizon at 23:30 (day 316) and since 1:21 UTC (day 317). No other known problems.
Effelsberg (A. Kraus at Wed Nov 13 08:56:27 2002) Success :
We chose another RX-filter at UT1445; that gives lower Tsys in channel 2.
Jodrell_Bank (Alastair Gunn at Fri Nov 15 13:46:53 2002) Minor failures : RFI
Missed first 8 minutes of schedule due to Rx repairs. Telescope
on source about 14:08 UT. Severe RFI at about 1320 MHz. No other
Westerbork (Tony Foley at Thu Dec 12 11:13:18 2002) Minor failures : HA-E
Scans after 01:35 out of HA range